Inversion of mappings Jan Malý Charles University in Prague

                    maly@karlin.mff.cuni.cz

Abstract: If f ∈ W^{1,n−1} is a homeomorphism, then the inverse mapping is in the BV class and the formula on differentiation of f^{−1} can be derived. For Sobolev regularity of the inverse mapping the assumption of finite distortion is relevant. Based on joint results with Katarína Bellová, Marianna Csörnyei, Stanislav Hencl, Pekka Koskela and Jani Onninen.